Solve the equation by a method of your choice.
4x^2+15x+9=0

Answer by rwm(914)

factor by grouping
what are factors of 36 (4*9) which add up to 15
12 and 3
4x^2+12x+3x+9=0
4x*(x+3)+3*(x+3)=0
(4x+3)*(x+3)=0
4x+3=0
x=-3/4
x+3=0
x=-3